# Affiliations withheld for double-blind review

### v2

@@ Abstract
Electronic health records (EHRs) enable powerful clinical prediction models but typically offer only coarse, post hoc explanations that are difficult to use in patient-level decision making. To address this gap, we propose a knowledge graph (KG)-guided chain-of-thought (CoT) framework that produces clinically grounded reasoning for visit-level disease prediction in MIMIC-III. We map ICD-9 codes to PrimeKG, mine disease-specific relevant nodes and reasoning paths, and use these KG paths as scaffolds to generate temporally consistent CoT explanations, retaining only samples whose conclusions match observed outcomes. We then fine-tune lightweight LLaMA-3.1-Instruct-8B and Gemma-7B models under two resulting training cohorts (400 and 1,000 index visits from MIMIC-III). On ten PrimeKG-mapped diseases and small training cohorts, our models outperform strong classical baselines, achieving AUROC of approximately 0.66–0.70 and macro-AUPR of 0.40–0.47. Models transfer zero-shot to the CRADLE cohort without any additional fine-tuning, improving accuracy from roughly 0.40–0.51 to 0.72–0.77, and blinded clinician evaluation shows a consistent preference for KG-guided CoT in terms of clarity, relevance, and correctness.
@@ Introduction
Electronic health records (EHRs) are now routinely collected across large and diverse healthcare systems, capturing longitudinal data on diagnoses, medications, and laboratory results from millions of patients.king2014clinical; ruan2019representation; zheng2014automatically As interest in digital medicine grows,fogel2018ai; landi2020deep these data are increasingly used to train predictive models for outcomes such as in-hospital mortality, readmission, and disease onset, often using benchmark cohorts like MIMIC-III.johnson2016mimic; goldstein2017practical Classical models such as logistic regression and tree-based ensembles remain popular due to their robustness, global interpretability, and ease of deployment in clinical systems,goldstein2017practical while neural architectures including recurrent and attention-based models have been explored to capture temporal dynamics and co-occurrence patterns in EHR data.choi2016retain; ruan2019representation Despite these advances in predictive performance, most EHR models provide only coarse and post hoc interpretability, offering limited support for patient-specific, clinician-facing reasoning about why a particular outcome is likely.
@@
Advances in large language models (LLMs) and reasoning-oriented architectures offer a potential remedy: LLMs can generate intermediate reasoning steps alongside predictions, enabling explicit chain-of-thought (CoT) narratives that articulate why a given prediction should hold.wei2022chain; wang2024a; huang202202cot; xie2024reasoning In principle, such step-by-step explanations could bridge the gap between accurate prediction and clinically usable justification for clinical predictions. However, applying CoT in medicine remains challenging. High-quality, clinically validated CoT datasets are scarce muennighoff2025; ye2025 and unconstrained LLM reasoning can introduce hallucinated or medically incorrect intermediate steps, which is particularly problematic in high-stakes clinical decision making.xu2024ram; liu2025ragreview These limitations have motivated efforts to constrain or ground LLM reasoning using external knowledge sources to improve factual reliability. Biomedical knowledge graphs (KGs) such as PrimeKG integrate curated relationships among diseases, phenotypes, drugs, and genes, and have been leveraged to enhance factual grounding in medical question answering and decision-support systems.chandak2023primekg; liu2025ragreview
@@
In this paper, we connect these previous works by integrating biomedical KG structure directly into the CoT generation process for visit-level disease prediction on structured EHR data. We propose a KG-guided CoT framework for MIMIC-IIIjohnson2016mimic that produces clinically grounded reasoning rather than unconstrained free-form explanations. Concretely, we map ICD-9 codes to PrimeKGchandak2023primekg through a three-stage entity-alignment procedure combining exact matching, embedding-based similarity retrieval, and LLM-driven clinical validation. Using these mapped entities, we identify disease-specific relevant features and extract KG paths linking each feature to its corresponding disease, pruning spurious or clinically irrelevant chains with GPT-4o. Guided by these paths, we prompt an LLM to generate temporally consistent, KG-anchored CoT explanations of whether a disease will appear at the next visit, and we retain only those reasoning traces whose conclusions match the ground-truth label. We then fine-tune lightweight open-weight models on this KG-anchored CoT supervision corpus so that, at inference time, they take ICD-9 features together with disease-specific KG evidence as input and produce both predictions and clinician-style explanations. This enables data-efficient, interpretable patient-level prognostic reasoning grounded in biomedical knowledge graphs.
@@
We evaluate this framework on ten PrimeKG-mapped diseases in MIMIC-III across small patient cohorts (400 and 1,000 patients), where our KG-anchored CoT supervision enables LLMs to outperform strong classical baselines such as XGBoost and Random Forest, achieving AUROC of approximately 0.67 and 0.70 and macro-AUPR of approximately 0.40 and 0.47, respectively. Clinician evaluations further show that KG-guided CoT yields substantially clearer, more coherent, and more clinically sound reasoning across three dimensions: clarity & coherence, coverage & relevance, and correctness & soundness. Finally, our CoT-tuned models transfer effectively to the CRADLE cohort for forecasting cardiovascular events within one year after type 2 diabetes, improving accuracy from roughly 0.40–0.51 (untuned LLMs) to 0.72–0.77 without any retraining, and providing more temporally consistent and clinically plausible explanations. Together, these results demonstrate that KG-anchored CoT supervision enables small LLMs to learn data-efficient, clinically grounded prognostic reasoning that generalizes across cohorts.
@@
A knowledge graph (KG) is a structured representation of entities and their relationships, typically modeled as a graph $G=(V,E)$ where $V$ denotes the set of entities (nodes) and $E$ the set of typed edges encoding relations among them. In the biomedical setting, KGs organize curated knowledge about diseases, phenotypes, drugs, genes, and related concepts, and provide an explicit relational backbone that can be used to guide or constrain model reasoning. Since biomedical relations in our external knowledge source are represented as direction- and type-specific (e.g., disease $\rightarrow$ gene and gene $\rightarrow$ disease may be encoded as distinct edges), we treat the underlying biomedical KG as a heterogeneous directed multigraph. Specifically, we adopt PrimeKGchandak2023primekg as our external knowledge source for grounding visit-level prognostic reasoning. PrimeKG is a large-scale biomedical KG that integrates curated relationships among diseases, phenotypes, drugs, genes, and other biomedical entities into a unified graph. We use it as the substrate for (1) aligning structured EHR concepts to KG nodes, (2) identifying disease-relevant neighborhood regions, and (3) extracting reasoning paths that connect clinical features to target diseases.
@@
Figure 1: General pipeline for KG-guided CoT data generation.
@@
where each $\hat{e}_{k,d}$ is a mapped feature node selected from $\{\hat{e}_{i}^{f}\}_{i=1}^{n}$ for disease $d$. Given $\mathcal{R}_{d}$, we extract reasoning chains from PrimeKG by computing all shortest paths between each feature node $\hat{e}_{k,d}\in\mathcal{R}_{d}$ and the disease node $\hat{e}_{d}$. Following prior work on KG-grounded reasoning luo2025o1pruner; chen2024overthinking, we use shortest paths to limit “overthinking” and preserve only the most immediate biomedical relations:
@@
Each generated CoT $C_{t,d}$ is required to end with an explicit binary conclusion $\hat{y}_{t+1}^{(d)}\in\{\text{Yes},\text{No}\}$. We use this conclusion to decide whether to keep the example: $C_{t,d}$ is kept if and only if the CoT-implied label matches the ground-truth outcome $y_{t+1}^{(d)}$, i.e., $\hat{y}_{t+1}^{(d)}=y_{t+1}^{(d)}$. The surviving set of $(x_{t},d,C_{t,d},y_{t+1}^{(d)})$ pairs forms a compact but high-quality KG-grounded supervision corpus. We fine-tune lightweight open-weight LLMs (e.g., LLaMA grattafiori2024llama, Gemma team2024gemma) on this dataset so that, at test time, they produce both predictions and clinician-style reasoning traces.
Figure 2: Schematic and example of KG-guided CoT generation and filtering.
@@ Results
We work with longitudinal visits from the MIMIC-III critical care database.johnson2016mimic Following the setup of Harutyunyan et al.,harutyunyan2019multitask we focus on patients with more than one hospital visit and construct all pairs of adjacent visits. For each such pair, the earlier visit serves as the index visit at time $t$ and the later visit provides labels for the next visit at time t+1t{+}1. This preprocessing yields 12,353 labeled index visits (i.e., visit pairs), which we refer to as patient cases in the remainder of the paper.
Each index visit at time $t$ is represented as a binary ICD-9 feature vector $x_{t}\in\{0,1\}^{N}$ with N=7,423N=7{,}423, where each dimension indicates the presence of a specific ICD-9 code during that visit. The original task is formulated as a multi-label prediction problem over 25 acute-care conditions derived from Clinical Classifications Software (CCS).harutyunyan2019multitask In this work, we adopt the same 25 disease labels but focus our evaluation on a subset $\mathcal{D}$ of $|\mathcal{D}|=10$ diseases that can be reliably mapped to PrimeKG (listed in Table 1). For each disease $d\in\mathcal{D}$ and index visit $t$, the prediction target is a binary label $y^{(d)}_{t+1}\in\{0,1\}$ indicating whether $d$ is present at the next visit t+1t{+}1. Models take as input $(x_{t},d)$ (and, for certain variants, KG evidence constructed as described in the Method section) and output a probability $\hat{p}^{(d)}_{t+1}$ and a binary verdict. We perform visit-level splitting over the 12,353 labeled index visits, randomly reserving 10% of cases as a held-out test set. From the remaining pool, we construct training sets of size 400 and 1,000 index visits to study data efficiency, using the rest for model selection.
In addition to MIMIC-III, we evaluate zero-shot transfer on Project CRADLE (Emory Clinical Research Analytics Data Lake Environment), a de-identified EHR repository from Emory Healthcare. Following prior work,xu2023hypehr we focus on patients with type 2 diabetes and formulate a binary prediction task: whether a patient will experience a
@@
As shown in Table 2, KG-guided CoT fine-tuning substantially boosts the performance of both LLaMA-3.1-8B and Gemma-7B in the low-data MIMIC-III setting. With only 400 labeled index visits data, the KG-CoT–tuned LLaMA model attains the best accuracy and macro-F1 and the highest macro-AUPR across all methods, while matching the AUROC of the strongest logistic regression baseline ($\approx 0.67$). Gemma-7B shows a similar trend: KG-CoT supervision raises its macro-AUPR from 0.25 to 0.40 and macro-F1 from essentially zero to $\approx 0.40$, bringing it in line with the tuned LLaMA variant. When the training set is increased to 1,000 cases, both KG-guided models improve further, with LLaMA-3.1 achieving the top accuracy, AUROC ($\approx 0.70$), and macro-AUPR ($\approx 0.47$), and Gemma-7B achieving the highest macro-F1, consistently outperforming strong classical baselines such as XGBoost and Random Forest. These results indicate that KG-anchored CoT supervision enables small open-weight LLMs to learn data-efficient, clinically grounded prognostic reasoning that is competitive with or superior to traditional EHR prediction models.
To evaluate cross-cohort robustness, we apply the MIMIC-trained models directly to the CRADLE dataset (36,611 patients, 12,724 binary features) without any re-training or calibration. The prediction task is cardiovascular disease onset within one year of type 2 diabetes diagnosis.
Table 3: Zero-shot transfer performance on the CRADLE cohort (cardiovascular disease within one year after T2D diagnosis). Best metric in bold, second-best underlined.
@@
MIMIC-trained models transfer effectively to the CRADLE cohort to forecast cardiovascular disease within one year after the diagnosis of type 2 diabetes. For the LLaMA-3.1 backbone, fine-tuning on 400–1,000 MIMIC-III cases increases accuracy from 0.40 to 0.75–0.77 and macro-F1 from 0.34 to 0.46–0.51, while modestly raising AUROC (to $\approx 0.52$–0.53) and AUPR (to $\approx 0.25$). Gemma-7B exhibits a comparable pattern: accuracy rises from 0.51 to 0.72–0.77 and macro-F1 from 0.31 to about 0.48, again with modest changes in AUROC and AUPR. These results suggest that the reasoning learned from KG-guided CoT supervision on MIMIC-III generalize across cohorts and feature spaces, enabling zero-shot transfer of both backbones to out-of-distribution EHR data.
@@
To evaluate whether KG-guided CoT improves explanation quality, we conducted a blinded human study comparing our KG-anchored LLaMA-3.1 model with its untuned baseline. For each of 115 randomly sampled cases, clinicians were shown an input summary, the ground-truth outcome, and two anonymized outputs (A and B), each containing a prediction and its reasoning trace, one from the KG-guided model and one from the baseline. Without knowing which model produced which trace, clinicians made pairwise preferences, selecting the better explanation (A or B) along three dimensions: 1. Clarity & Coherence: whether the explanation is easy to follow, with logically ordered steps and clear transitions. 2. Coverage & Relevance: whether the explanation identifies key factors, focus on clinically relevant information, and avoids unnecessary detail. 3. Correctness & Soundness: whether the clinical statements, risk factors, and temporal relationships are accurate and logically justified, without internal contradictions.
For each dimension, annotators were instructed to choose whichever output (A or B) they judged to be better. The annotation was performed by two human experts from different medical schools: an MD professor and a MD/PhD student in biomedical informatics. In total, 115 unique cases were annotated, with the MD professor reviewing 65 cases and the MD/PhD student reviewing 50 cases.
Figure 3: Clinician preferences for reasoning quality.
On the 115 annotated cases, both clinicians preferred the KG-guided model (Model A) over the untuned baseline (Model B), with overall 96.5% of cases for Clarity & Coherence, 94.8% of cases for Coverage & Relevance, and 98.3% of cases for Correctness & Soundness. Qualitatively, clinicians reported that KG-guided CoT traces more often (i) organized evidence in a temporally consistent manner, (ii) highlighted clinically meaningful risk factors rather than superficial code patterns, and (iii) avoided unsupported leaps or hallucinated details. By contrast, the untuned baseline was described as more “trigger-happy” in predicting positive outcomes (more likely to conclude YES even when only weak signals were present). While both models are sensitive to poor or incomplete input structure, the untuned baseline was more prone to repetitive differentials and illogical reasoning, especially on shorter or noisier inputs.

### v3

@@ Abstract
Electronic health records (EHRs) support strong clinical prediction but often provide coarse, post hoc explanations that are hard to use for patient-level decisions. We propose a knowledge-graph (KG)–guided chain-of-thought (CoT) framework for visit-level disease prediction on MIMIC-III. We map ICD-9 codes to PrimeKG, mine disease-relevant nodes and paths, and use these paths to scaffold temporally consistent CoT explanations, retaining only samples whose conclusions match observed outcomes. We then fine-tune lightweight LLaMA-3.1-Instruct-8B and Gemma-7B models on two small cohorts (400 and 1,000 index visits) across ten PrimeKG-mapped diseases. Our models outperform strong classical baselines, reaching AUROC of 0.66–0.70 and macro-AUPR of 0.40–0.47. Without additional training, the models transfer zero-shot to the CRADLE cohort, improving accuracy from 0.40–0.51 to 0.72–0.77. Blinded clinicians consistently prefer KG-guided CoT for clarity, relevance, and correctness. Code is available at https://github.com/JonathanWry/KG-guided-LLM-pipeline.
@@ Introduction
Electronic health records (EHRs) are now routinely collected across large and diverse healthcare systems, capturing longitudinal data on diagnoses, medications, and laboratory results from millions of patients.11; 16; 25 As interest in digital medicine grows,5; 12 these data are increasingly used to train predictive models for outcomes such as in-hospital mortality, readmission, and disease onset, often using benchmark cohorts like MIMIC-III.10; 6 Classical models such as logistic regression and tree-based ensembles remain popular due to their robustness, global interpretability, and ease of deployment in clinical systems,6 while neural architectures including recurrent and attention-based models have been explored to capture temporal dynamics and co-occurrence patterns in EHR data.4; 16 Despite these advances in predictive performance, most EHR models provide only coarse, post hoc interpretability (such as global feature importance, attribution scores, or attention weights), offering limited support for patient-specific, clinician-facing reasoning about why a particular outcome is likely.6; 4; 1; 17
@@
Advances in large language models (LLMs) and reasoning-oriented architectures offer a potential remedy: LLMs can generate intermediate reasoning steps alongside predictions, enabling explicit chain-of-thought (CoT) narratives that articulate why a given prediction should hold.20; 19; 9; 21 In principle, such step-by-step explanations could bridge the gap between accurate prediction and clinically usable justification for clinical predictions. However, applying CoT in medicine remains challenging. High-quality, clinically validated CoT datasets are scarce 15; 24 and unconstrained LLM reasoning can introduce hallucinated or medically incorrect intermediate steps, which is particularly problematic in high-stakes clinical decision making.23; 13 These limitations have motivated efforts to constrain or ground LLM reasoning using external knowledge sources to improve factual reliability. Biomedical knowledge graphs (KGs) such as PrimeKG integrate curated relationships among diseases, phenotypes, drugs, and genes, and have been leveraged to enhance factual grounding in medical question answering and decision-support systems.2; 13
@@
In this paper, we connect these previous works by integrating biomedical KG structure directly into the CoT generation process for visit-level disease prediction on structured EHR data. We propose a KG-guided CoT framework for MIMIC-III10 that produces clinically grounded reasoning rather than unconstrained free-form explanations. In our formulation, each visit-level instance uses ICD-9 code features from an index visit $t$ to predict a binary disease label at the immediately subsequent visit t+1t{+}1. Concretely, we map ICD-9 codes to PrimeKG2 through a three-stage entity-alignment procedure combining exact matching, embedding-based similarity retrieval, and LLM-driven clinical validation. Using these mapped entities, we identify disease-specific relevant features and extract KG paths linking each feature to its corresponding disease, pruning spurious or clinically irrelevant chains with GPT-4o. Guided by these paths, we prompt an LLM to generate temporally consistent, KG-anchored CoT explanations of whether a disease will appear at the next visit, and we retain only those reasoning traces whose conclusions match the ground-truth label. We then fine-tune lightweight open-weight models on this KG-anchored CoT supervision corpus so that, at inference time, they take ICD-9 features together with disease-specific KG evidence as input and produce both predictions and clinician-style explanations. This enables data-efficient, interpretable patient-level prognostic reasoning grounded in biomedical knowledge graphs.
@@
We evaluate this framework on ten PrimeKG-mapped diseases in MIMIC-III across small patient cohorts (400 and 1,000 patient index visits), where our KG-anchored CoT supervision enables LLMs to outperform strong classical baselines such as XGBoost and Random Forest, achieving AUROC of approximately 0.67 and 0.70 and macro-AUPR of approximately 0.40 and 0.47, respectively. Clinician evaluations further show that KG-guided CoT yields substantially clearer, more coherent, and more clinically sound reasoning across three dimensions: clarity & coherence, coverage & relevance, and correctness & soundness. Finally, our CoT-tuned models transfer effectively to the CRADLE cohort for forecasting cardiovascular events within one year after type 2 diabetes, improving accuracy from roughly 0.40–0.51 (untuned LLMs) to 0.72–0.77 without any retraining, and providing more temporally consistent and clinically plausible explanations. Together, these results demonstrate that KG-anchored CoT supervision enables small LLMs to learn data-efficient, clinically grounded prognostic reasoning that generalizes across cohorts.
@@
Figure 1: General pipeline for KG-guided CoT data generation. Observed clinical features are denoted in Blue; Intermediate KG nodes are denoted in Green, and the Target disease is denoted in Red.
@@
A knowledge graph (KG) is a structured representation of entities and their relationships, typically modeled as a graph $G=(V,E)$ where $V$ denotes the set of entities (nodes) and $E$ the set of typed edges encoding relations among them. In the biomedical setting, KGs organize curated knowledge about diseases, phenotypes, drugs, genes, and related concepts, and provide an explicit relational backbone that can be used to guide or constrain model reasoning. Since biomedical relations in our external knowledge source are represented as direction- and type-specific (e.g., disease $\rightarrow$ gene and gene $\rightarrow$ disease may be encoded as distinct edges), we treat the underlying biomedical KG as a heterogeneous directed multigraph. Specifically, we adopt PrimeKG2 as our external knowledge source for grounding visit-level prognostic reasoning. PrimeKG is a large-scale biomedical KG that integrates curated relationships among diseases, phenotypes, drugs, genes, and other biomedical entities into a unified graph. We use it as the substrate for (1) aligning structured EHR concepts to KG nodes, (2) identifying disease-relevant neighborhood regions, and (3) extracting reasoning paths that connect clinical features to target diseases.
@@
where each $\hat{e}_{k,d}$ is a mapped feature node selected from $\{\hat{e}_{i}^{f}\}_{i=1}^{n}$ for disease $d$. Given $\mathcal{R}_{d}$, we extract reasoning chains from PrimeKG by computing all shortest paths between each feature node $\hat{e}_{k,d}\in\mathcal{R}_{d}$ and the disease node $\hat{e}_{d}$. Following prior work on KG-grounded reasoning 14; 3, we use shortest paths to limit “overthinking” and preserve only the most immediate biomedical relations:
@@
Each generated CoT $C_{t,d}$ is required to end with an explicit binary conclusion $\hat{y}_{t+1}^{(d)}\in\{\text{Yes},\text{No}\}$. We use this conclusion to decide whether to keep the example: $C_{t,d}$ is kept if and only if the CoT-implied label matches the ground-truth outcome $y_{t+1}^{(d)}$, i.e., $\hat{y}_{t+1}^{(d)}=y_{t+1}^{(d)}$. The surviving set of $(x_{t},d,C_{t,d},y_{t+1}^{(d)})$ pairs forms a compact but high-quality KG-grounded supervision corpus. We fine-tune lightweight open-weight LLMs (e.g., LLaMA 7, Gemma 18) on this dataset so that, at test time, they produce both predictions and clinician-style reasoning traces.
Figure 2: Schematic and example of KG-guided CoT generation and filtering. Observed clinical features are denoted in Blue; Intermediate KG nodes are denoted in Green, and the Target disease is denoted in Red.
@@ Results
We work with longitudinal visits from the MIMIC-III critical care database.10 Following the setup of Harutyunyan et al.,8 we focus on patients with more than one hospital visit and construct all pairs of adjacent visits. For each such pair, the earlier visit serves as the index visit at time $t$ and the later visit provides labels for the next visit at time t+1t{+}1. This preprocessing yields 12,353 labeled index visits (i.e., visit pairs).
Each index visit at time $t$ is represented as a binary ICD-9 feature vector $x_{t}\in\{0,1\}^{N}$ with N=7,423N=7{,}423, where each dimension indicates the presence of a specific ICD-9 code during that visit. The original task is formulated as a multi-label prediction problem over 25 acute-care conditions derived from Clinical Classifications Software (CCS).8 In this work, we adopt the same 25 disease labels but evaluate on |D|=10|D|{=}10 diseases that map cleanly to PrimeKG (Table 1). For each disease $d\in\mathcal{D}$ and index visit $t$, the prediction target is a binary label $y^{(d)}_{t+1}\in\{0,1\}$ indicating whether $d$ is present at the next visit t+1t{+}1. Models take as input $(x_{t},d)$ (and, for certain variants, KG evidence constructed as described in the Method section) and output a probability $\hat{p}^{(d)}_{t+1}$ and a binary verdict. e split 12,353 index visits by visit (10% test) and subsample training sets of 400 and 1,000 visits for data-efficiency experiments, using the remainder for validation.
In addition to MIMIC-III, we evaluate zero-shot transfer on Project CRADLE (Emory Clinical Research Analytics Data Lake Environment), a de-identified EHR repository from Emory Healthcare. Following prior work,22 we focus on patients with type 2 diabetes and formulate a binary prediction task: whether a patient will experience a
@@
As shown in Table 2, KG-guided CoT fine-tuning substantially boosts the performance of both LLaMA-3.1-8B and Gemma-7B in the low-data MIMIC-III setting. With only 400 labeled index visits data, the KG-CoT–tuned LLaMA model attains the best accuracy and macro-F1 and the highest macro-AUPR across all methods, while matching the AUROC of the strongest logistic regression baseline ($\approx 0.67$). Gemma-7B shows a similar trend: KG-CoT supervision raises its macro-AUPR from 0.25 to 0.40 and macro-F1 from essentially zero to $\approx 0.40$, bringing it in line with the tuned LLaMA variant. When the training set is increased to 1,000 cases, both KG-guided models improve further, with LLaMA-3.1 achieving the top accuracy, AUROC ($\approx 0.70$), and macro-AUPR ($\approx 0.47$), and Gemma-7B achieving the highest macro-F1, consistently outperforming classical baselines such as XGBoost and Random Forest. These results indicate that KG-anchored CoT supervision enables small open-weight LLMs to learn data-efficient, clinically grounded prognostic reasoning that is competitive with or superior to traditional EHR prediction models.
Table 3: Ablation on MIMIC-III visit-level disease prediction with 1000 patient index visits. Best metric in bold, second-best underlined.
[…]
LLaMA3-8B + KG (unfiltered) 0.8485 0.6589 0.4473 0.4023
LLaMA3-8B (no KG, filtered) 0.8613 0.6640 0.4637 0.4357
[…]
Table 4: Zero-shot transfer on the CRADLE cohort. Best metric in bold, second-best underlined.
@@
To isolate KG guidance and filtering effects, we compare three variants at 1,000 index visits in Table 3. Removing KG guidance or using unfiltered CoT traces reduces AUROC/AUPR, suggesting that ungrounded or noisy rationales degrade ranking performance. Overall, KG-grounded and curated CoT supervision is key to strong AUROC/AUPR in this regime.
To assess cross-cohort robustness, we evaluate the MIMIC-trained models zero-shot on CRADLE (36,611 patients; 12,724 binary features) to predict cardiovascular disease within one year after type 2 diabetes diagnosis, without re-training or calibration. MIMIC-trained models transfer to the CRADLE cohort to forecast cardiovascular disease within one year after the diagnosis of type 2 diabetes. For the LLaMA-3.1 backbone, fine-tuning on 400–1,000 MIMIC-III cases increases accuracy from 0.40 to 0.75–0.77 and macro-F1 from 0.34 to 0.46–0.51, while modestly raising AUROC (to $\approx 0.52$–0.53) and AUPR (to $\approx 0.25$). Gemma-7B exhibits a comparable pattern: accuracy rises from 0.51 to 0.72–0.77 and macro-F1 from 0.31 to about 0.48, again with modest changes in AUROC and AUPR. These results suggest that the reasoning learned from KG-guided CoT supervision on MIMIC-III generalize across cohorts and spaces, enabling zero-shot transfer of backbones to out-of-distribution EHR data.
@@
To evaluate whether KG-guided CoT improves explanation quality, we conducted a blinded human study comparing our KG-anchored LLaMA-3.1 model with its untuned baseline. For each of 115 randomly sampled cases, clinicians were shown an input summary, the ground-truth outcome, and two anonymized model outputs (A and B), each containing a prediction and its reasoning trace (one KG-guided, one baseline). Without knowing which model produced which trace, clinicians made a pairwise preference judgment (A vs. B) along three complementary dimensions:
  1. 1.
    Clarity & Coherence (presentation): he trace is easy to follow, logically ordered, and avoids jumps, circular reasoning, or contradictions.
  2. 2.
    Coverage & Relevance (evidence selection): the trace surfaces key patient-specific factors for the target disease, stays clinically focused, and avoids generic or distracting content.
  3. 3.
    Correctness & Soundness (clinical validity): clinical claims and timelines are medically plausible and consistent with the provided features/outcome (no fabricated details or inappropriate causal leaps), and the conclusion follows from the stated evidence.
These dimensions are intentionally distinct: clarity evaluates how the reasoning is communicated, coverage evaluates what evidence is selected and emphasized, and correctness evaluates whether the medical content is accurate and logically warranted. The annotation was performed by two human experts from different medical schools: an MD professor and an MD/PhD student in biomedical informatics. In total, 115 unique cases were annotated, with the MD professor reviewing 65 cases and the MD/PhD student reviewing 50 cases.
Inter-rater consistency. To quantify consistency across adjudicators, we additionally double-annotated a random subset of $n_{\text{overlap}}$ of 55 cases (both clinicians independently rated the same cases across all three dimensions) and computed Cohen’s $\kappa$ on the binary preference decision (KG-guided vs. baseline). Agreement was $\kappa_{\text{clar}}$ for Clarity & Coherence, $\kappa_{\text{cov}}$ for Coverage & Relevance, and $\kappa_{\text{corr}}$ for Correctness & Soundness.
Figure 3: Clinician preference for KG-guided explanations vs. the untuned baseline across three dimensions (donuts show % preferred; remainder is baseline).
On the 115 annotated cases, both clinicians preferred the KG-guided model (Model A) over the untuned baseline (Model B): 96.5% for Clarity & Coherence, 94.8% for Coverage & Relevance, and 98.3% for Correctness & Soundness. Qualitatively, clinicians noted that KG-guided CoT traces more often (i) organized evidence temporally, (ii) highlighted clinically meaningful risk factors over superficial code patterns, and (iii) avoided unsupported leaps or hallucinated details. By contrast, the untuned baseline was described as more “trigger-happy” (more likely to conclude YES from weak signals). While both models are sensitive to poor or incomplete inputs, the baseline was more prone to repetitive differentials and illogical reasoning, especially on short or noisy inputs.